OpenStudy (anonymous):

Can someone please help me. I need to see how to work this. I will give a medal. Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33

OpenStudy (cggurumanjunath):

f(x)=x+8 f(2)=2+8

OpenStudy (anonymous):

i think you need g(2) first right?

OpenStudy (cggurumanjunath):

g(x) = x2 - 6x - 7 g(2)=(2)^2-6(2)-7

OpenStudy (anonymous):

first find g(2) then take f(result)

OpenStudy (cggurumanjunath):

g(2)=(2)^2-6(2)-7=4-12-7=-8-7=-15

OpenStudy (cggurumanjunath):

f(g(2))=f(-15)

OpenStudy (cggurumanjunath):

f(x)=x+8 f(-15)=-15+8

OpenStudy (anonymous):

g(2) = (2)^2-6(2)-7 = -15 f(15) = -15 +8 is your answer

OpenStudy (anonymous):

ok thank you both for explaining it, I appreciate it. have a great day.

OpenStudy (cggurumanjunath):

how much is -15+8 that is your final answer @mlb004

OpenStudy (cggurumanjunath):

welcome.